# latexText page

### v1

@@ 5.3 Relationship between word importance and concreteness
The relationship between average word importance and the word frequency in the training corpus is plotted in Fig. 16, with different sample colors representing the levels of concreteness. There is no obvious relationship between the word attention and the frequency. Many high frequency words obtain a low word importance when the concreteness value is very low, and on the opposite, low frequency words of concrete concepts may obtain a high word importance. Fig. 16 (right) shows the zoom-in on the samples in the blue box, which have high average word importance and in the purple box with low average importance. Comparing these two boxes, the words with concrete visual meaning, especially nouns, are indeed more likely to obtain higher importance than the abstract words. Therefore, the analysis using Grad-ECLIP text explanation further supports the conclusion from §5.2concrete visual concepts are learned better and contribute more to the CLIP image-text matching score than abstract concepts, and this phenomenon is not due to word frequency bias.
6 Application of Grad-ECLIP to boost fine-grained understanding of CLIP
CLIP has been shown to have limitations in understanding fine-grained details, such as poor region recognition when using its dense features, due to the pre-training focusing on matching the whole image (via the $[cls]$ token) to a text description. Our Grad-ECLIP generates text-specific explanation maps for the image encoder, which highlights the detailed regions on the input image when CLIP matches the image-text pair. Therefore, aiming to boost the fine-grained understanding ability of CLIP, here we propose a fine-tuning method for CLIP that adopts Grad-ECLIP to indicate the fine-grained alignment between image regions and corresponding textual attributes. As a result, the fine-tuned CLIP has a representation space where visual and semantic features are both globally and locally aligned, which significantly improves its region aware ability for dense prediction tasks, as well as maintaining its performance on image-level tasks.
Figure 17: Overview of the proposed fine-grained fine-tuning of CLIP using Grad-ECLIP. Multiple phrases or words representing objects (e.g., “dog”, “black car” and “traffic lights”) are separated out by parsing the input caption via the Natural Language Toolkit (NLTK). With the visual explanation heat maps generated by Grad-ECLIP, object-specific region feature embeddings ($F_{r}$) are obtained through weighted aggregation of the image dense feature ($F_{d}$). Finally, the CLIP fine-tuning is composed of two losses: the global loss as in the original pre-training of CLIP, and the local loss, which aligns the image region features ($F_{r}$) and the corresponding phrase features ($F_{p}$).
6.1 Grad-ECLIP-based fine-grained fine-tuning of CLIP
The framework of our proposed method is shown in Fig. 17, where the fine-tuned model includes an image encoder $\mathcal{I}\left(\cdot\right)$ and a text encoder $\mathcal{T}\left(\cdot\right)$, which is the same model architecture as the original CLIP. In the next two subsections, we will introduce the global loss and local loss adopted in the fine-grained fine-tuning, respectively.
6.1.1 Global loss
The input of CLIP is a batch of image-text pairs $\{(I_{b},T_{b})\}_{b=1}^{B}$. After passing through the image encoder and text encoder, the model outputs the corresponding global image and text embeddings $\{(F_{I_{b}},F_{T_{b}})\}_{b=1}^{B}$, respectively. The fine-tuning adopts the same global contrastive learning as in the pre-training, which realizes the instances-level alignment and helps to maintain the model’s ability for multi-modal global representation. The cosine similarity $S(F_{I},F_{T})$ between image embedding $F_{I}$ and text embedding $F_{T}$ is calculated as in Eq. 1. The constrastive loss is applied to push CLIP to learn the global representations by maximizing the cosine similarities of the corresponding image and text embeddings, while minimizing the cosine similarities of other non-paired ones in the batch, which is defined as:
$L_{global}=-\dfrac{1}{2B}\sum_{b=1}^{B}\left(\log\tfrac{\exp\left(S(F_{I_{b}},F_{T_{b}})/\tau\right)}{\sum_{b^{\prime}=1}^{B}\exp\left(S(F_{I_{b}},F_{T_{b^{\prime}}})/\tau\right)}+\right.$ (21)
$\left.\log\tfrac{\exp\left(S(F_{T_{b}},F_{I_{b}})/\tau\right)}{\sum_{b^{\prime}=1}^{B}\exp\left(S(F_{T_{b}},F_{I_{b^{\prime}}})/\tau\right)}\right),$
where $\tau$ is the trainable temperature parameter.
6.1.2 Local loss
The local loss is based on the extracting the dense feature map from the image and performing fine-grained matching of region features to the corresponding text descriptions.
Image dense feature. Following [19, 28] we extract the dense feature map of the input image from a ViT-based encoder by to slightly modifying the last transformer layer to keep the projection and norm layers, and discard the self-attention. This modification is experimentally shown to be capable of preserving more spatial detailed features in the output token embeddings. Specifically, in the last transformer layer, with the input $x=(x_{cls},x_{1},...,x_{h\times w})$ comprising a $[cls]$ embedding and $h\times w$ spatial token embeddings, the output of the attention layer is $o=v=\mathcal{LP}(x)$ instead of the $o=\mathcal{A}(x)=\mathrm{softmax}(\tfrac{qk^{\mathsf{T}}}{\sqrt{C}})v$ in (7). Then, the $[cls]$ embedding is removed and the final spatial token embeddings are reshaped into an $h\times w$ image dense feature map $F_{d}$, from which we extract fine-grained representations for specific image regions.
Image-text fine-grained matching. Next we design a region feature and text matching scheme based on Grad-ECLIP to automatically obtain region-text alignment pairs. Note that our method does not require any manual or network-based region proposal or label annotations. For the caption $T$ in each image-text pair $(I,T)$, we use the Natural Language Toolkit (NLTK) [82] to parse and extract the phrases that contain object concepts, by setting the separation and selection rules as “adjective + noun”. For the example in Fig. 17, the NLTK extracts “dog”, “black car”, and “traffic lights” from the input text “a dog in a black car waiting for traffic lights”. Then, these extracted words or phrases $\{p_{t}\}_{t=1}^{n}$, where $n$ is the maximum number of extracted concepts from each caption, are sent to the text encoder, resulting in a set of phrase embeddings $\{F_{p_{t}}\}_{t}$. A region embedding $F_{r_{t}}$ is then calculated for each phrase embedding. Specifically, the phrase embeddings $F_{p_{t}}$ is used to calculate the cosine similarity with the image embedding $F_{I}$. Grad-ECLIP is then applied to the calculated score to obtain a heat map $H_{t}$ according to the procedure in §3.2, which reveals the important spatial locations for matching with this specific phrase. Therefore, we adopt the explanation heat maps as weights for aggregating the image dense feature $F_{d}$, resulting in the region embedding for the phrase: $F_{r_{t}}=\sum_{hw}H_{t}\cdot F_{d}$, where $\sum_{hw}$ is the sum operator over spatial coordinates.
Finally, for the region embeddings and corresponding phrase embeddings, we adopt focal loss [64] to match the positive pairs and distance the negative pairs:
$L_{local}=-\sum_{t}\left(1-S\left(F_{r_{t}},F_{p_{t}}\right)\right)^{2}\log S\left(F_{r_{t}},F_{p_{t}}\right)$ (22)
$-\sum_{t}\sum_{t^{\prime}\neq t}S\left(F_{r_{t}},F_{p_{t^{\prime}}}\right)^{2}\log\left(1-S\left(F_{r_{t}},F_{p_{t^{\prime}}}\right)\right),$
where $S$ represents the cosine similarity function, and $t$, $t^{\prime}$ means the $t$-th and $t^{\prime}$-th phrase in the same batch. By adding the local loss to the global loss as the total loss, our proposed fine-grained fine-tuning successfully boosts the representation alignment between image region and corresponding textual concepts, while maintains the image-level performance at the same time.
6.2 Experiments with fine-grained fine-tuning CLIP
In these experiments, we use our proposed Grad-ECLIP to enable fine-tuning of CLIP to enhance the fine-grained understanding.
6.2.1 Experiment settings
The experiments are conducted based on the pre-trained models from EVA-CLIP [83] considering its high efficiency and capacity, followed the previous work [28]. For the fine-tuning experiments, unless otherwise specified, we adopt the Conceptual Caption (CC3M) dataset [71] as the training set, which collects about 3 million image-text pairs from the internet. During fine-tuning, the input image size is 224x224, which is the same as the original pre-training of CLIP. Two RTX 6000 Ada are used with batch size 64 on each, learning rate of 1e-5, and weight decay of 0.1.
6.2.2 Experiment results
TABLE VII: Evaluating the fine-grained representation of the fine-tuned CLIP via zero-shot classification on the MS COCO validation dataset. We report the Top-1 and Top-5 mean accuracy on both object bounding boxes and panoptic masks (thing and stuff). Besides our Grad-ECLIP, we also adopt Grad-CAM and MaskCLIP to produce explanation map in the calculation of local loss for comparison. The red arrow shows the performance improvement brought by the local loss based on our Grad-ECLIP, compared with just using the global loss, which equivalent to ordinary fine-tuning. The gray row is the baseline CLIP before fine-tuning.
Fine-tuning Boxes Thing Masks Stuff Masks
Method Model Global Loss Local Loss Explanation Map Top1 Top5 Top1 Top5 Top1 Top5
CLIP ViT-B/16 - - - 41.4 63.6 30.6 53.8 13.9 36.6
ViT-B/16 $\surd$ - - 42.9 64.8 32.9 56.4 14.7 38.7
Fine-tuned $\surd$ $\surd$ Grad-CAM 54.2 74.7 46.5 69.8 13.2 42.2
CLIP $\surd$ $\surd$ MaskCLIP 54.3 75.5 47.4 70.9 17.0 47.9
$\surd$ $\surd$ Grad-ECLIP (Ours) 57.3↑14.4 78.3↑13.5 49.3↑16.4 72.2↑15.8 18.3↑3.6 51.1↑12.4
CLIP ViT-L/14 - - - 58.1 78.9 49.8 72.6 13.1 33.9
Fine-tuned ViT-L/14 $\surd$ - - 62.6 83.1 54.7 77.5 16.2 39.2
CLIP $\surd$ $\surd$ Grad-ECLIP (Ours) 71.7↑9.1 89.7↑6.6 63.4↑8.7 85.6↑8.1 20.5↑4.3 53.6↑14.4
Evaluating the fine-grained representation. To evaluate the dense representation ability of the fine-tuned CLIP, we use the mean accuracy (mAcc) of classifying region boxes annotated in the val2017 split of MS COCO and panoptic masks (including “things” with 80 classes and “stuffs” with 91 classes) annotated in MS COCO Panoptic dataset [84]. For classification, RoI pooling is used to extract region box embeddings, while mask pooling is used to extract the mask embeddings from the image dense feature maps. The classification is performed by selecting the highest score when matching with the text embeddings of the classes.
The results are shown in Tab. VII. Compared with the pre-trained CLIP base model, fine-tuning with just the global loss on CC3M dataset can slightly increase the zero-shot classification performance – this setting is equivalent to continuing to train CLIP in the original way, which we denote as ordinary FT (fine-tuning) in the following descriptions. When we use both the global and local loss with the help of our Grad-ECLIP explanation maps, the fine-tuned model’s performances on region classification obtained significant improvements, for both boxes and masks. In Tab. VII, we also list the absolute increases of each metric compared with ordinary FT (global loss only) using the red values, which shows the effectiveness of the local loss on boosting the fine-grained representation of CLIP ViT model.
To demonstrate the effectiveness of our Grad-ECLIP, we also conduct the fine-tuning with the local loss using two other plug-in and low computation cost explanation methods, Grad-CAM and MaskCLIP. With the heat maps from these two methods, the region-aware matching also obtains obvious performance improvements compared with the ordinary FT, but there is still a gap compared with using our Grad-ECLIP. The comparison results further demonstrate that the high-quality and accuracy of the visual explanation maps generated by Grad-ECLIP.
TABLE VIII: Results on open-vocabulary object detection on MS COCO val set. F-ViT is the two-stage detector baseline built on the frozen original CLIP ViT, and $\dagger$ means the ViT backbone is initialized with the CLIP model fine-tuned (FT) with the corresponding method on the dataset in brackets (CC3M or MS COCO Karpathy trainset). Ordinary FT is equivalent to just using the global loss in the fine-tuning.
[…]
OV-RCNN [85] ResNet50 17.5 41.0 34.9
RegionCLIP [13] ResNet50 26.8 54.8 47.5
Detic [86] ResNet50 27.8 51.1 45.0
VLDet [87] ResNet50 32.0 50.6 45.8
F-VLM [56] ResNet50 28.0 - 39.6
CORA [57] ResNet50 35.1 35.5 35.4
RO-ViT [27] ViT-B/16 30.2 - 41.5
RO-ViT [27] ViT-L/16 33.0 - 47.7
F-ViT ViT-B/16 19.4 43.3 37.0
+CLIPSelf [28] (CC3M) ViT-B/16 13.4 39.3 32.5
+Ordinary FT (CC3M) ViT-B/16 19.5 43.4 37.1
+Our FT (CC3M) ViT-B/16 27.4↑8.0 43.8↑0.5 39.5↑2.5
+CLIPSelf [28] (MS COCO) ViT-B/16 25.2 42.2 37.7
+Ordinary FT (MS COCO) ViT-B/16 20.1 43.8 37.6
+Our FT (MS COCO) ViT-B/16 26.7↑7.3 44.2↑0.9 39.6↑2.6
F-ViT ViT-L/14 28.3 52.5 46.2
+Ordinary FT (CC3M) ViT-L/14 31.1 53.4 47.6
+Our FT (CC3M) ViT-L/14 39.4↑11.1 53.6↑1.1 49.9↑3.7
Application to open-vocabulary detection (OVD) task. We adopt the fine-tuned CLIP as the backbone for OVD to verify the fine-grained understanding by a down-stream localization task. Following the previous work CLIPSelf [28], which is a state-of-the-art CLIP fine-tuning scheme for increasing the fine-grained region representation via self-distillation, we build open-vocabulary object detectors based on the F-ViT [28] architecture, which is a two-stage detector using a frozen CLIP ViT as the backbone. In the CLIP fine-tuning stage, we adopt the image-text pairs in CC3M or MS COCO Karpathy train set as the inputs, and for MS COCO Karpathy train set, we filter out the same image samples that are also in the OVD val set to ensure there is no risk of label leakage. The OVD models are trained on the OV-COCO benchmark [88], and we use AdamW optimizer with batch size of 64, learning rate of 1e-4, and weight decay of 0.1. For evaluation, we report box AP (average precision) at IoU (Intersection over Union) of base, novel and all categories as with previous works [13, 56, 57, 27, 28]. Since there is no extra region box annotations used in our fine-tuning, for fair comparison, we implement fine-tuning of the CLIPSelf version using image patch distillation, which also has no extra region proposal annotation requirements.
The results are presented in Tab. VIII. F-ViT is the baseline that initializes the detector backbone with the original pre-trained CLIP model, and other versions with $\dagger$ initialize the backbone with the CLIP models fine-tuned by various methods and datasets. With the ViT-B/16 backbone, ordinary FT produces similar performance as the baseline, while our FT significantly improves the OVD results, especially on the novel categories. Since the base categories have explicit annotated bounding boxes and labels during OVD training, the performance on the unseen novel categories better illustrates the fine-grained understanding ability brought by the CLIP model. In contrast to CLIPSelf where the downstream OVD performance is much influenced by the dataset used for fine-tuning, our method obtains better performances regardless of the fine-tuning dataset (CC3M or MS COCO), which indicates that our method can effectively and stably boost the dense representation ability of the CLIP model. Finally, we conduct the experiments with the ViT-L/14 architecture and further improve the OVD performance on the novel categories by a large extent. Compared with the existing OVD methods, mostly relying on ResNet-based encoder or modified ViT encoder, and requiring pre-training from scratch on prepared large-scale data with extra region information, our method achieves superior performance with just fine-tuning the CLIP on the easily-obtained image-text pairs.
TABLE IX: Evaluating the influence of fine-grained fine-tuning on image-level representation by a zero-shot retrieval task using Flicker30k. R@i denotes the recall accuracy with top $i$ matching. CLIP is the pre-trained CLIP model from EVA-CLIP [83], which is the base model the fine-tuning (FT) methods adopt. Ordinary FT is equivalent to just using the global loss in the fine-tuning.
text-to-image image-to-text
Method Model R@1 R@5 R@10 R@1 R@5 R@10
CLIP [83] ViT-B/16 73.6 90.9 94.8 88.6 97.1 99.1
+CLIPSelf [28] ViT-B/16 46.9 72.8 81.1 51.4 76.9 85.5
+Ordinary FT ViT-B/16 72.7 90.7 94.5 86.7 96.7 98.5
+Our FT ViT-B/16 72.8 91.0 94.7 88.3 97.4 98.9
CLIP [83] ViT-L/14 78.8 93.9 96.8 90.4 98.8 99.4
+Ordinary FT ViT-L/14 80.2 94.5 97.1 91.6 98.9 99.7
+Our FT ViT-L/14 80.2 94.6 97.1 92.1 99.1 99.8
Evaluating the influence on image-level representation. We next explore the influence of our fine-tuning on the image-level representation ability, to see if there are negative effects to image-level representation when improving fine-grained representations. We report the recall accuracy of image-to-text and text-to-image retrieval task with the Flickr30k [33] validation set. As the results shown in Tab. IX, after fine-tuning the CLIP ViT-B/16 model on the CC3M dataset, our FT has successfully preserve the image-level retrieval performance, which is similar to the ordinary FT. In contrast, CLIPSelf has largely lost the ability of its global image-level representation. Finally, the superior results on ViT-L/14 model compared with the baseline CLIP further demonstrate the effectiveness of our FT in maintaining both the global representation and stable image-level matching.
7 Conclusion
In this paper, we propose Grad-ECLIP, a novel white-box gradient-based visual and textual explanation method for CLIP, the dual-encoder pre-trained model for image-text matching. Grad-ECLIP is applicable to both the image and text encoders, producing heat maps that indicate the importance of image regions or words for the image-text matching score. Qualitative and quantitative evaluations demonstrate the advantages of Grad-ECLIP compared with existing explanation methods designed for transformers/CLIP, and the adaptation experiments exhibit the generalizability of our method. We also adopt Grad-ECLIP to analyze the properties of the pre-trained CLIP model, where we discover its ability of concept decomposition and addibility, advantages/limitations on different attribute identification, and its bias towards learning concrete words over abstract words. By introducing these analyses as examples, we hope the proposed interpretation method can be used to help with both development and understanding of VLMs. Finally, we propose a fine-tuning scheme, which adopts the Grad-ECLIP explanation map to obtain region-text pairs for a local loss that boosts the fine-grained understanding of CLIP. In future work, we will consider how to associate individual words from the sentence to regions in the image, and vice versa. Future work can also consider how to extend Grad-ECLIP to other VLMs with modified dual-encoder architectures, e.g ALBEF with additional cross-attention layers to fuse image and text features.
Acknowledgements
This work was supported by Strategic Research Grants from City University of Hong Kong (Project. Nos. 7005840 and 7005995) and Hetao Shenzhen-Hong Kong Science and Technology Innovation Cooperation Zone (HZQB-KCZYZ-2021045).

### v2

@@ V-C Relationship between word importance and concreteness
The relationship between average word importance and the word frequency in the training corpus is plotted in Fig. 17, with different sample colors representing the levels of concreteness. There is no obvious relationship between the word attention and the frequency. Many high-frequency words obtain a low word importance when the concreteness value is very low, and on the opposite, low-frequency words of concrete concepts may obtain a high word importance. Fig. 17 (right) shows the zoom-in on the samples in the blue box, which have high average word importance, and in the purple box with low average importance. Comparing these two boxes, the words with concrete visual meaning, especially nouns, are indeed more likely to obtain higher importance than the abstract words. Therefore, the analysis using Grad-ECLIP text explanation further supports the conclusion from §V-Bconcrete visual concepts are learned better and contribute more to the CLIP image-text matching score than abstract concepts, and this phenomenon is not due to word frequency bias.
V-D Limitations and future work
In this paper, we have analyzed CLIP via XAI methods by observing and interpreting the explanation heat maps on different samples, such as in Sections V-I and V-II. However, it should be noted that there are limitations of heat map-based analysis, which stem from two aspects. Firstly, the observation results are illustrated through a small set of examples (compared to the size of the dataset), where limited samples will reduce the credibility of conclusions. Secondly, the faithfulness of any XAI method cannot reach a hundred percent, and the importance ranks obtained by certain methods cannot perfectly display both the necessity and sufficiency [60]. Therefore, conclusions drawn from XAI results from certain methods are subject to some extent. In future work, we will consider when given image-sentence pair (with a sentence containing multiple words), how to associate each important word from the sentence to each important region in the image, and vice versa. Future work can also consider how to extend Grad-ECLIP to other VLMs with modified dual-encoder architectures, e.g ALBEF with additional cross-attention layers to fuse image and text features.
VI Conclusion
In this paper, we propose Grad-ECLIP, a novel white-box gradient-based visual and textual explanation method for CLIP, the dual-encoder pre-trained model for image-text matching. Grad-ECLIP is applicable to both the image and text encoders, producing heat maps that indicate the importance of image regions or words for the image-text matching score. Qualitative and quantitative evaluations demonstrate the advantages of Grad-ECLIP compared with existing explanation methods designed for transformers/CLIP, and the adaptation experiments exhibit the generalizability of our method. Finally, we also adopt Grad-ECLIP to analyze the properties of the pre-trained CLIP model, where we discover its ability of concept decomposition and addibility, advantages/limitations on different attribute identification, and its bias towards learning concrete words over abstract words. By introducing these analyses as examples, we hope the proposed interpretation method can be used to help with both development and understanding of VLMs.
Acknowledgments
This work was supported by Strategic Research Grants from City University of Hong Kong (Project. Nos. 7030010 and 7005995).